# OccluTrack: Rethinking Awareness of Occlusion for Enhancing Multiple Pedestrian Tracking

### v1

@@ IV-D Ablation Study
TABLE I: The results of components ablation study on three different existing trackers over MOT17 validation set. MS KF means our abnormal motion suppression Kalman Filter. PG re-ID stands for pose-guided re-ID. OD Mea. is our proposed occlusion-aware distance measurement.
Tracker MS KF PG re-ID OD Mea. IDF1 (%) ↑ IDSw↓
@@
We investigate the impact of alternative speed filters on our tracker’s performance, which are the Gaussian filter, Laplacian filter, and Mean Average filter. Each filter is applied to smooth the buffered centering moving and scale-changing speeds. The Gaussian filter convolves the buffered speeds with a Gaussian kernel, effectively smoothing the centering moving speed and scale-changing speed. The Laplacian filter employs a weighted average of neighboring speeds to smooth the buffered speeds based on their distances. Finally, the Mean Average filter computes the average of all the buffered speeds, as described in Section III-B. We summarize the comparisons among these filters based on the MOT17 validation set in Table II. The results in Table II indicate that our proposed Mean average filter yields the best performance in terms of IDF1 and IDSw. This filter achieves the highest accuracy in estimating pedestrian moving speeds, leading to improved tracking performance.
@@
We analyzed the impact of different feature embedding fusion methods on the performance of OccluTrack. The evaluated fusion techniques included addition, concatenation, and our proposed adaptive fusion. The results on the MOT17 validation set, presented in Table III, indicate that the adaptive fusion method outperformed the addition and concatenation methods, demonstrating its effectiveness in improving tracking performance.
@@
Figure 5: Illustration of the IDF1 and MOTA during adjusting abnormal motion suppression parameter $\alpha_{0}$.
In this section, we ablate the abnormal motion suppression parameter $\alpha_{0}$ on the MOT17 validation set. We vary $\alpha_{0}$ from 0 to 1 with step 0.1 and evaluate its impact on the tracking performance. $\alpha_{0}=0$ means the Kalman Filter does not update when abnormal motions are detected, while $\alpha_{0}=1$ means the Kalman Filter does not suppress abnormal motions and operates as the normal Kalman Filter. The results are illustrated in Fig.5. When increasing $\alpha_{0}$ from 0 to 0.2, we observe that the IDF1 reached the highest at $\alpha_{0}=0.2$ and MOTA increased from the lowest 76.4% at $\alpha_{0}=0$ to the highest 77.3% at $\alpha_{0}=0.2$. Further increasing $\alpha_{0}$ leads to a gradual decline in IDF1 and reaches its lowest 81.5% at $\alpha_{0}=1.0$, while MOTA remains relatively stable and ranges between 77.2% and 77.3%. Based on these findings, we set $\alpha_{0}=0.2$ for evaluating the MOT17 and MOT20 test sets.
@@ IV-E Benchmark Evaluation
TABLE IV: Evaluation results and comparisons with the state-of-the-art-trackers on MOT17 test set under private detection protocols.
Tracker IDF1↑ HOTA↑ MOTA↑ FP↓ FN↓ IDSw↓ AssA↑ AssR↑
STTD [29] 51.9 - 54.2 11435 244443 2735 - -
QuasiDense [27] 66.3 53.9 68.7 26589 146643 3378 52.7 57.2
FairMOT [8] 72.3 59.3 73.7 27507 117477 3303 58.0 63.6
TransCenter [36] 62.2 54.5 73.2 23112 123738 4614 49.7 54.2
TransTrack [13] 63.5 54.1 75.2 50157 86442 3603 47.9 57.1
CSTrack [52] 72.6 59.3 74.9 23847 114303 3567 57.9 63.2
RelationTrack [53] 74.7 61.0 73.8 27999 118623 1374 61.5 67.3
TransMOT [14] 75.1 61.7 76.7 36231 93150 2346 59.9 66.5
FDTrack [33] 75.6 61.3 76.8 34467 92718 3705 - -
MOTFR [32] 76.3 61.8 74.4 32397 109245 2625 62.6 67.8
OCSORT [22] 77.5 63.2 78.0 15129 107055 1950 63.2 67.5
StrongSORT++ [34] 79.5 64.4 79.6 27876 86205 1194 64.4 71.0
ByteTrack [9] 77.3 63.1 80.3 25491 83721 2196 62.0 68.2
Quo Vadis [19] 77.7 63.1 80.3 - - 2103 62.1 68.8
BoT-SORT [10] 80.2 65.0 80.5 22521 86037 1212 65.5 71.2
Ours 81.1 65.4 80.6 22311 85971 1038 66.3 71.8
TABLE V: Evaluation results and comparisons with the state-of-the-art-trackers on MOT20 test set under private detection protocols.
Tracker IDF1↑ HOTA↑ MOTA↑ FP↓ FN↓ IDSw↓ AssA↑ AssR↑
SiamMOT [54] 69.1 - 67.1 - - - - -
FairMOT [8] 67.3 54.6 61.8 103440 88901 5243 54.9 60.7
TransCenter [36] 50.4 - 61.9 45895 146347 4653 37.0 45.1
TransTrack [13] 59.4 48.5 65.0 27197 150197 3608 45.2 51.9
CSTrack [52] 68.6 54.0 66.6 25404 144358 3196 54.0 57.6
RelationTrack [53] 70.5 56.5 67.2 61134 104597 4243 56.4 60.3
SOTMOT [55] 71.4 - 68.6 57064 101154 4209 - -
MOTFR [32] 71.7 57.2 69.0 54894 101823 3648 57.1 62.6
FDTrack [33] 75.7 59.9 75.0 24011 102896 2226 - -
OCSORT [22] 76.3 62.4 75.7 19067 105894 942 62.5 67.4
StrongSORT++ [34] 77.0 62.6 73.8 16632 117920 770 64.0 69.6
ByteTrack [9] 75.2 61.3 77.8 26249 87594 1223 59.6 66.2
Quo Vadis [19] 75.7 61.5 77.8 - - 1185 59.9 67.0
BoT-SORT [10] 77.5 63.3 77.8 24638 88863 1257 62.9 68.6
Ours 78.6 64.1 77.9 25079 87943 1124 65.6 70.3
We compare our OccluTrack with the state-of-the-art trackers on MOT17 and MOT20 datasets under private detection protocols, as shown in Table IV and Table V. Our OccluTrack consistently outperforms other state-of-the-art trackers across multiple important metrics.
@@
On the MOT17 dataset, our OccluTrack excels all other trackers in terms of five metrics, including IDF1, HOTA, MOTA, AssA, and AssR. Compared to our baseline tracker, BoT-SORT, our method consistently outperforms it on all metrics and achieves state-of-the-art performance. Regarding IDF1 and MOTA, our method improves 0.9% and 0.1% compared with our baseline, which indicates that our method has a more powerful tracking performance while keeping compatible detection performance. Correspondingly, the HOTA improved by 0.5%, considering both detection and tracking performance. Noticeably, the IDSw decreases by 14.2% from 1212 and reaches the optimal performance at 1038. Our method also surpasses all the other methods on AssA and AssR, showing more promising association ability. Although our method cannot achieve state-of-the-art performance on metrics like FP and FN compared with OC-SORT [22] and ByteTrack [9], it still outperforms our baseline.
@@
Even though MOT20 is a more complex dataset with more crowdedness compared with the MOT17 dataset, our OccluTrack still achieves promising performance. Especially our proposed OccluTrack achieves better performance regarding IDF1, HOTA, MOTA, AssA, and AssR. For IDF1 and MOTA, OccluTrack improves them by 1.1% and 0.1 %, corresponding to tracking performance and detection performance, contributing to 0.8% improvements on HOTA. This demonstrates that OccluTrack is superior in tracking while keeping good detection performance. In terms of association performance, our OccluTrack shows the best performance and achieves dramatic improvements, as high as 2.7% and 1.7% on AssA and AssR. While OccluTrack is not as good as OC-SORT [22] and ByteTrack [9] on FP, FN, and IDSw, it still performs better than the baseline. Particularly for IDSw, our OccluTrack outperforms our baseline and reduces it by 10.6% from 1257 to 1124, showing more stable tracking performance.
Figure 6: Comparisons of BoT-SORT and our proposed OccluTrack on six samples from MOT17 and MOT20 datasets. On MOT17, we visualize sampled persons and corresponding trajectories to show the tracking performance. On MOT20, we show the sampled frames with selected video since it is hard to show the trajectories because of the crowdedness.
@@ IV-F Visualization
We conducted a comprehensive analysis of tracking results obtained from six videos in the MOT17 validation and test sets, as shown in Fig.6. Our purpose is to compare the tracking performance of our OccluTrack with that of the current state-of-the-art BoT-SORT under various conditions. We show the results of the two trackers when the selected pedestrians are before partial occlusion, during partial occlusion, during occlusion, and after occlusion. Besides, we demonstrate the trajectories for both trackers on the right of the figure.
@@
As shown in Fig.6, BoT-SORT struggled to assign the correct ID to the person with ID 39, whose ID changed to 47 after occlusion. In contrast, our OccluTrack consistently tracked the same person with ID 38 throughout the sequence. This discrepancy in performance can be attributed to the superior accuracy of OccluTrack’s predictions during occlusion, as they are closely aligned with the positions of occluded pedestrians. The visualized trajectories further confirm the robust tracking capability of our OccluTrack in maintaining track continuity for the person with ID 38. Similar trends were observed from other sub-figures, where our OccluTrack successfully tracked the person with ID 34 in MOT17-02-FRCNN, ID 30 in MOT17-06-FRCNN, ID 9 in both MOT17-08-FRCNN and MOT17-09-FRCNN, and ID 16 in MOT17-10-FRCNN, but BoT-SORT encountered difficulties.
@@
In the MOT20 dataset, due to crowded scenes that limit visualization, we selected samples from all two videos in the test set to showcase the tracking performance of our OccluTrack. We sample three frames from each video at intervals of 200 frames and present the final tracking results in Fig.6. Due to the challenges posed by crowded scenes, it was not feasible to visualize individual trajectories or specific tracks. However, the visualization, combined with the corresponding numerical results in Section IV-E, shows the impressive tracking performance of our OccluTrack in handling crowded pedestrian tracking within the MOT20 dataset.
Our comprehensive evaluation of tracking results over MOT17 and MOT20 datasets demonstrates the superior tracking performance of our OccluTrack compared to SOTA. The visualizations provide evidence of OccluTrack’s effectiveness in tackling occlusion challenges and accurately tracking pedestrians in different scenarios.
V Conclusion
In conclusion, we have demonstrated that abnormal motion caused by partial occlusion is the missing key to enhancing multiple pedestrian tracking performance in IDF1 and IDSw. Our proposed OccluTrack can alleviate occlusion problems in multiple pedestrian tracking. By introducing an abnormal motion suppression mechanism, a pose-guided re-ID module, and an occlusion-aware distance measurement, we effectively mitigated the negative impact of partial occlusion on motion estimation, appearance features, and associations. Extensive evaluations showed OccluTrack outperforms state-of-the-art methods, particularly in IDF1, IDSw, HOTA, AssA, and AssR metrics.

### v2

@@ IV-D Ablation Study
TABLE I: The results of the components ablation study on three different existing trackers over the MOT17 validation set. We applied these three modules (AMS KF, PG Re-ID, and OD Mea.) to three different existing trackers (JDE, FairMOT, and BoT-SORT) to demonstrate the effectiveness of our proposed method.
Tracker AMS KF PG Re-ID OD Mea. IDF1 (%) ↑ IDSw↓
@@
We investigate the impact of alternative speed filters, the Gaussian filter, the Laplacian filter, and the Mean Average filter, on our OccluTrack. Each filter is applied to smooth the buffered centering, moving, and scale-changing speeds. The Gaussian filter convolves the buffered speeds with a Gaussian kernel, effectively smoothing the centering, moving speed, and scale-changing speed. The Laplacian filter employs a weighted average of neighboring speeds to smooth the buffered speeds based on their distances. Finally, the Mean Average filter computes the average of all the buffered speeds, as described in Section III-B. We summarize the comparisons among these filters based on the MOT17 validation set in Table II. The results in Table II indicate that our proposed Mean average filter yields the best performance in terms of IDF1 and IDSw. This filter achieves the highest accuracy in estimating pedestrian moving speeds, leading to improved tracking performance.
@@
We analyzed the impact of different feature embedding fusion methods on the performance of OccluTrack. The evaluated fusion techniques included addition, concatenation, and our proposed adaptive fusion. The results on the MOT17 validation set, presented in Table III, indicate that the adaptive fusion method outperforms the addition and concatenation methods, demonstrating its effectiveness in improving tracking performance.
@@
Figure 5: Ablation study of abnormal motion suppression parameter $\alpha_{0}$ on the MOT17 validation set. The $\alpha_{0}$ is adjusted from 0 to 1. IDF1 and MOTA jointly reach the peak when setting $\alpha_{0}$=0.2.
Figure 6: Ablation study of abnormal motion detection threshold $\theta_{v}$ on the MOT17 validation set. The $\theta_{v}$ is adjusted from 0 to 1. The best IDF1 and MOTA are jointly obtained when setting $\theta_{v}$=0.2.
In this section, we ablate the abnormal motion suppression parameter $\alpha_{0}$ on the MOT17 validation set. We vary $\alpha_{0}$ from 0 to 1 with a step of 0.1 and evaluate its impact on the tracking performance. $\alpha_{0}=0$ means the Kalman Filter does not update when abnormal motions are detected, while $\alpha_{0}=1$ means the Kalman Filter does not suppress abnormal motions and operates as the normal Kalman Filter. The results are illustrated in Fig. 6. When increasing $\alpha_{0}$ from 0 to 0.2, we observe that the IDF1 reached the highest at $\alpha_{0}=0.2$ and MOTA increased from the lowest 76.4% at $\alpha_{0}=0$ to the highest 77.3% at $\alpha_{0}=0.2$. Further increasing $\alpha_{0}$ leads to a gradual decline in IDF1 and reaches its lowest 81.5% at $\alpha_{0}=1.0$, while MOTA remains relatively stable and ranges between 77.2% and 77.3%. Based on these findings, we set $\alpha_{0}=0.2$ for evaluating the MOT17 and MOT20 test sets.
Abnormal Motion Detection Threshold
Apart from the suppression parameter, we also conduct ablation studies on the effect of adjusting the abnormal motion detection threshold $\theta_{v}$ from 0 to 1 with a step of 0.1. The results of IDF1 and MOTA are illustrated in Fig. 6. When increasing $\theta_{v}$ from 0 to 0.2, both MOTA and IDF1 achieve the optimal points 77.3 % and 83.0 %. With further increasing the threshold, the joint performance of MOTA and IDF1 fluctuates and drops to 77.1 % and 81.3 %. In summary, the optimal threshold for abnormal motion detection should be set to 0.2 to obtain the best tracking performance.
@@ IV-E Benchmark Evaluation
We compare our OccluTrack with the state-of-the-art trackers on MOT17 [21], MOT20 [22], and DanceTrack [23] datasets. We provide two versions of our proposed tracker, OccluTrack and OccluTrack+, building on BoT-SORT [13] and BoostTrack++ [57], respectively. For MOT17, we benchmark our OccluTrack and OccluTrack+ under private detection protocol and OccluTrack under public protocol in Table IV and Table V. For the private protocol, the detection results are from our object detector, which is trained on multiple datasets. For the public protocol, we filter the YOLOX detection results by calculating the Intersection over Union (IoU) with the detection results provided by MOTChallenge, following ByteTrack [12]. These results come from detectors such as FRCNN, SDP, and DPM, and we apply an IoU threshold of 0.5 for filtering. For MOT20, we only compare OccluTrack and OccluTrack+ with the state-of-the-art method under the private detection protocols in Table VI as limited published methods provide the results under the public detection protocol. We also provide benchmark comparisons on the DanceTrack dataset under the private detection protocol.
@@
Under the private protocol, our OccluTrack excels BoT-SORT on all metrics. Regarding IDF1 and MOTA, our method improves by 1.8% and 1% compared with our baseline, which indicates that our method has a more powerful tracking performance while keeping compatible detection performance. Correspondingly, the HOTA improved by 0.7%, considering both detection and tracking performance. Compared with the most recent work, our OccluTrack outperforms UCMCTrack on IDF1, MOTA, and IDSw by 1.0%, 0.9%, and 19.0%. As for the OccluTrack+, it outperforms existing methods on most metrics like IDF1, HOTA, and IDSw, achieving state-of-the-art performance. Compared with BoostTrack++ (baseline), our OccluTrack+ improves IDF1 and HOTA to 82.8% (0.6% ↑) and 66.8% (0.2% ↑) while decreasing the ID Switches to 951 (10.5% ↑). The improvements demonstrating our OccluTrack+ show better tracking performance by mitigating the partial occlusion.
Under public detection, our proposed OccluTrack outperforms existing state-of-the-art methods that provide results under public detection by a large margin. As the results in Table V indicate, our methods achieved the best performance among the existing methods on all listed important metrics. Specifically, OccluTrack improves the IDF1, HOTA, MOTA, and IDSw by 7.6%, 5.2%, 5.8%, and 35.3% compared with ByteTrack [12]. The improvements show that our OccluTrack has superior tracking performance among existing methods. Noticeably, the proposed tracker can be modified and plugged into any tracking-by-detection method.
@@
Even though MOT20 is a more complex dataset with more crowdedness compared with the MOT17 dataset, our OccluTrack and OccluTrack+ still achieve promising performance. Comparing our OccluTrack with BoT-SORT, it achieves better performance in IDF1, HOTA, MOTA, and IDSw. For IDF1 and MOTA, OccluTrack improves them by 1.2% and 0.1%, corresponding to tracking performance and detection performance, contributing to 0.8% improvements on HOTA. This demonstrates that OccluTrack is superior in tracking while keeping good detection performance. As for IDSw, our OccluTrack outperforms our baseline and reduces it by 10.6% from 1257 to 1124, showing a more stable tracking performance. As for OccluTrack+, it achieves state-of-the-art performance regarding IDF1 and HOTA, improving BoostTrack++ by 0.7% and 0.3%, which also decreases IDSw. from 762 to 429. It is evident that our proposed methods are capable of improving tracking performance by considering partial occlusion, especially decreasing the ID Switches in crowded scenes.
DanceTrack
Apart from MOT17 and MOT20 datasets, we also evaluate our proposed OccluTrack on the DanceTrack dataset to further evaluate the effectiveness of fast-motion scenarios. Different from MOT17 and MOT20, we set the abnormal motion suppression parameters $\alpha_{0}$ and $\theta_{v}$ to 0.7, which is less sensitive to fast motion changes. As the results in Table. VII show, we benchmark our OccluTrack and methods used in MOT17 and MOT20 for fair comparisons. We focus on the three most important and mostly reported metrics: IDF1, HOTA, and MOTA. Our proposed OccluTrack also outperforms existing methods, especially OC-SORT [25], with noticeable improvements. For IDF1, HOTA, and MOTA, OccluTrack surpasses the OC-SORT [25] by 3.7%, 1.8%, and 0.2% respectively. The improvements of IDF1 and HOTA also demonstrate advanced tracking performance over OC-SORT.
TABLE IV: Evaluation results and comparisons with the state-of-the-art-trackers on MOT17 test set under the private detection protocol.
Tracker IDF1↑ HOTA↑ MOTA↑ IDSw↓
STTD [30] 51.9 - 54.2 2735
QuasiDense [29] 66.3 53.9 68.7 3378
FairMOT [11] 72.3 59.3 73.7 3303
TransCenter [58] 62.2 54.5 73.2 4614
TransTrack [15] 63.5 54.1 75.2 3603
CSTrack [59] 72.6 59.3 74.9 3567
RelationTrack [60] 74.7 61.0 73.8 1374
TransMOT [16] 75.1 61.7 76.7 2346
FDTrack [61] 75.6 61.3 76.8 3705
MOTFR [62] 76.3 61.8 74.4 2625
OCSORT [25] 77.5 63.2 78.0 1950
StrongSORT++ [33] 79.5 64.4 79.6 1194
ByteTrack [12] 77.3 63.1 80.3 2196
Quo Vadis [19] 77.7 63.1 80.3 2103
BoT-SORT [13] 80.2 65.0 80.5 1212
UCMCTrack [63] 81.0 65.7 80.6 1689
BoostTrack+ [57] 81.8 66.4 80.6 1086
BoostTrack++ [57] 82.2 66.6 80.7 1062
Ours (OccluTrack) 82.0 65.7 81.5 1368
Ours (OccluTrack+) 82.8 66.8 80.2 951
TABLE V: Evaluation results and comparisons with the state-of-the-art-trackers on MOT17 test set under the public detection protocol.
Tracker IDF1↑ HOTA↑ MOTA↑ IDSw↓
Tracktor [64] 55.1 44.8 56.3 1985
CenterTrack [26] 59.6 48.2 61.5 2583
TMOH [65] 62.8 50.4 62.1 1897
QuasiDense [29] 65.1 - 64.6 2652
SiamMOT [66] 63.3 - 65.9 3040
Trackformer [14] 57.6 - 62.3 4018
METFormer [34] 59.4 - 62.9 2827
ByteTrack [12] 70.0 56.1 67.4 1331
Ours (OccluTrack) 77.6 61.3 73.2 984
TABLE VI: Evaluation results and comparisons with the state-of-the-art trackers on MOT20 test set under the private detection protocol.
Tracker IDF1↑ HOTA↑ MOTA↑ IDSw↓
SiamMOT [67] 69.1 - 67.1 -
FairMOT [11] 67.3 54.6 61.8 5243
TransCenter [58] 50.4 - 61.9 5653
TransTrack [15] 59.4 48.5 65.0 3608
CSTrack [59] 68.6 54.0 66.6 3196
RelationTrack [60] 70.5 56.5 67.2 4243
SOTMOT [66] 71.4 - 68.6 4209
MOTFR [62] 71.7 57.2 69.0 3648
FDTrack [61] 75.7 59.9 75.0 2226
OCSORT [25] 76.3 62.4 75.7 942
StrongSORT++ [33] 77.0 62.6 73.8 770
ByteTrack [12] 75.2 61.3 77.8 1223
Quo Vadis [19] 75.7 61.5 77.8 1185
BoT-SORT [13] 77.5 63.3 77.8 1257
UCMCTrack [63] 77.4 62.8 75.6 1370
BoostTrack+ [57] 81.5 66.2 77.2 827
BoostTrack++ [57] 82.0 66.4 77.7 762
Ours (OccluTrack) 78.6 64.1 77.9 1124
Ours (OccluTrack+) 82.7 66.7 77.7 429
TABLE VII: Evaluation results and comparisons with the state-of-the-art trackers on DanceTrack test set.
Tracker IDF1↑ HOTA↑ MOTA↑
CenterTrack [26] 35.7 41.8 86.8
FairMOT [11] 40.8 39.7 82.2
QuasiDense [29] 44.8 45.7 83.0
TransTrack [15] 45.2 45.5 88.4
SORT [10] 50.8 47.9 91.8
DeepSORT [32] 47.9 45.6 87.8
ByteTrack [12] 52.5 47.3 89.5
OC-SORT [25] 54.9 55.1 92.2
Ours (OccluTrack) 59.6 56.9 92.4
@@ IV-F Visualization
We conducted a comprehensive analysis of tracking results obtained from two videos in the MOT17 and one video from the MOT20, as shown in Fig. 7.
@@
On the MOT17 dataset, our purpose is to compare the tracking performance of our OccluTrack with that of the state-of-the-art BoT-SORT under various conditions. We show the results of the two trackers when the selected pedestrians are before partial occlusion, during partial occlusion, during occlusion, and after occlusion. Besides, we demonstrate the trajectories for both trackers on the right of the figure with the occlusion period highlighted. As shown in Fig. 7 of MOT17-01, BoT-SORT struggled to assign the correct ID to the person with ID 39, whose ID changed to 47 after occlusion. In contrast, our OccluTrack consistently tracked the same person with ID 38 throughout the sequence. This discrepancy in performance can be attributed to the superior accuracy of OccluTrack’s predictions during occlusion, as they are closely aligned with the positions of occluded pedestrians. The visualized trajectories further confirm the robust tracking capability of our OccluTrack in maintaining track continuity for the person with ID 38. From the zoomed occlusion period, we can see that BoT-SORT is sensitive to abnormal motions caused by partial occlusion, while our OccluTrack is more robust when occlusion happens. A similar situation happens in MOT17-02. After occlusion, our OccluTrack consistently tracks the person (ID 34) without being affected by partial occlusion, while BoT-SORT fails to track the same person (ID 35).
@@
On the MOT20 dataset, due to crowded scenes that limit visualization, we selected two examples (ID 24 and 72) in MOT20-04 to showcase the tracking performance of our OccluTrack. We sample three frames from the video at intervals of 200 frames and present the final tracking results in Fig. 7. Due to the challenges posed by crowded scenes, it was not feasible to visualize individual trajectories or specific tracks. However, the visualization of example pedestrians (ID 24 and 72), combined with the corresponding numerical results in Section IV-E, shows the impressive tracking performance of our OccluTrack in handling crowded pedestrian tracking within the MOT20 dataset.
Figure 7: Visualization of our proposed OccluTrack on samples from the MOT17 and MOT20 datasets. On MOT17, we visualize sampled persons and corresponding trajectories with the occlusion period highlighted to show the tracking performance. In MOT-01, our OccluTrack consistently tracks the person with ID 38, while BoT-SORT fails to track the same person with ID 39. A similar situation happens to MOT17-02 for the person with ID 34 in OccluTrack and ID 35 in BoT-SORT.On MOT20, we use the sampled frames in MOT20-04 with two examples (ID 72 and 24) to show that our OccluTrack is also able to handle multiple pedestrian tracking in crowded scenarios.
Figure 8: Two examples of typical failure cases. In the first example, when a small-scale pedestrian (ID 31) is occluded in a long period captured by a fast-moving camera, our OccluTrack fails to track the pedestrian and assigns the wrong ID with 46 once it reappears. In the second example, the false negative pedestrian with ID 84 (not exist in the video sequence) is tracked due to the wrong detection results.
V Limitation and Future Work
Inference Speed with Re-ID. The most noticeable limitation of OccluTrack is the slight inference speed drop when compared with our baseline BoT-SORT. Based on the same object detector, the main effects come from two perspectives: pose estimation and the number of pedestrians. For the pose estimation, we compare our OccluTrack with pose-guided Re-ID with the BoT-SORT without pose guidance. With a pose estimator introduced, our current implementation operates at about 7.5 frames per second (FPS) on the MOT17 validation set, while BoT-SORT runs at 8 FPS. However, when the Re-ID module is removed, we can achieve a faster inference speed comparable to methods like BoT-SORT, reaching up to 30 FPS with a 0.4% drop on IDF1 evaluated on the MOT17 validation set. For the number of pedestrians, we compare the FPS and performance when applying OccluTrack to the MOT17 and MOT20 test sets. MOT17 is constructed with an average pedestrian density of 32, while MOT20 possesses an average pedestrian density of 171. As for the inference speed, the inference speed declined from 7.5 to 1.6 when changing MOT17 to MOT20. We also tested our OccluTrack without pose-guided Re-ID, and the inference speed increased to 7.5 with a tracking performance drop of 0.8 regarding IDF1.
Failure Cases. As shown in Fig. 8, we highlight two typical failure cases. The first occurs when a small-scale pedestrian is occluded for an extended period while captured by a fast-moving camera. In such cases, the tracking algorithm often struggles to maintain identity due to the challenges posed by the rapid camera movement. For example, the pedestrian with ID 31 in the first example of Fig. 8, its ID changes to 46 after such a situation. Specifically, the small scale of the pedestrian, combined with the fast motion of the camera, can cause significant deviations in the pedestrian’s trajectory. To mitigate this issue, more advanced techniques, such as optical flow, should be utilized to provide better motion compensation for the fast-moving camera. The second failure case arises from false negatives produced by the object detector. If the detector gives the wrong detections like ID 84 (not exist in the video sequence) in the second example of Fig. 8, the tracking algorithm assigns the false negative track. To address this, it is crucial to explore more robust object detectors that can minimize such errors.
VI Conclusion
In conclusion, we have shown that abnormal motion caused by partial occlusion is the missing key to enhancing multiple pedestrian tracking performance, and our proposed OccluTrack effectively addresses partial occlusion challenges through three key modules: a plug-and-play abnormal motion suppression mechanism within the Kalman Filter to detect and suppress outlier motions, a pose-guided Re-ID module that captures part-level features of partially visible pedestrians, and an occlusion-aware association strategy for associating occluded pedestrians with a fair distance measurement. Extensive evaluations of MOTChallenge and DanceTrack demonstrate that both OccluTrack and OccluTrack+ outperform the state-of-the-art methods from quantitative and qualitative results. For future work, we aim to optimize the inference speed in crowded scenarios, especially considering the computational overhead introduced by the Re-ID module. In addition, we will explore robust motion modeling and object detection techniques to address failure cases caused by fast-moving cameras and false negative detections.